OpenStudy (anonymous):

Help, pretty please? Which value of y forms the solution of the system defined by y = 11 x + 1 and 11 x + 12 y = 12?

jimthompson5910 (jim_thompson5910):

Start with 11x + 12y = 12 and plug in y = 11x + 1 11x + 12y = 12 11x + 12(11x + 1) = 12 ... replace every y with 11x + 1 now solve for x

OpenStudy (anonymous):

Ahhhhh,it's one of those problem,s..

jimthompson5910 (jim_thompson5910):

substitution is one way to do it, yes

jimthompson5910 (jim_thompson5910):

elimination is another (you have to rearrange things though)

OpenStudy (anonymous):

I got x=0 :/

OpenStudy (anonymous):

143x+12=12 x=0...

jimthompson5910 (jim_thompson5910):

good, x = 0 is correct

jimthompson5910 (jim_thompson5910):

now use this to find y

jimthompson5910 (jim_thompson5910):

you can plug x = 0 into any equation with x,y in it y = 11x + 1 is the easiest to work with since y is already isolated

jimthompson5910 (jim_thompson5910):

y = 11x + 1 y = 11(0) + 1 ... replace every x with 0 y = ??

OpenStudy (anonymous):

y=1

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

together, the solution is (0,1) ie x = 0, y = 1

OpenStudy (anonymous):

much appreciated!

jimthompson5910 (jim_thompson5910):

you're welcome